OpenStudy (anonymous):

generate the first five terms in the sequence using the explicit formula. yn=-5n-5

OpenStudy (anonymous):

plug in n = 1 for the first term, n =2 for the second term n = 3 for the third term and so on

OpenStudy (confluxepic):

1. Generate the first five terms in the sequence using the explicit formula. yn = –5n – 5 (1 point) (0 pts) –30, –25, –20, –15, –10 (0 pts) 30, 25, 20, 15, 10 (1 pt) –10, –15, –20, –25, –30 (0 pts) 10, 15, 20, 25, 30 1 /1 point 2. What is the 15th term in the sequence using the given formula? cn = 3n – 1 (1 point) (0 pts) 14 (0 pts) 57 (1 pt) 44 (0 pts) –44 1 /1 point 3. Write a recursive formula for the sequence 7, 13, 19, 25, 31, ... Then find the next term. (1 point) (1 pt) an = an– 1 + 6, where a1 = 7; 37 (0 pts) an = an– 1 + 6, where a1 = 37; 7 (0 pts) an = an– 1 – 6, where a1 = 6; –23 (0 pts) an = an– 1 – 6, where a1 = 7; –8 1 /1 point 4. Write a recursive formula for the sequence 7, 4, 1, –2, –5, .... Then find the next term. (1 point) (0 pts) an = an–1 – 3, where a1 = –8; 7 (1 pt) an = an–1 – 3, where a1 = 7; –8 (0 pts) an = an–1 + 3, where a1 = –3; 22 (0 pts) an = an–1 + 3, where a1 = 7; –8 1 /1 point 5. Write an explicit formula for the sequence 8, 6, 4, 2, 0, ... Then find a14. (1 point) (0 pts) an = –2n + 10; –16 (0 pts) an = –2n + 8; –18 (0 pts) an = –2n + 8; –20 (1 pt) an = –2n + 10; –18 1 /1 point

OpenStudy (confluxepic):

@~ @confluxepic